OpenStudy (anonymous):

Derive the equation of the parabola with a focus at (6, 2) and a directrix of y = 1

OpenStudy (anonymous):

f(x) = −one half (x − 6)2 + three halves f(x) = one half (x − 6)2 + three halves f(x) = −one half (x + three halves)2 + 6 f(x) = one half (x + three halves)2 + 6